Q.2  If following are the demand and supply equations for a consumer good:   (20)

        Qd = 300,000/P (Qd = quantity demanded)

        Qs = 30P (Qs = quantity supplied)

        (P = Price)

 

           Find the market clearing price and the quantity at the equilibrium. Also sketch the demand and supply curves.


1
Expert's answer
2022-01-07T09:25:30-0500

At Equilibrium, "Q^d=Q^s"

"\therefore \frac{300000}{P}=30P"

"30P^2=300000"

"P^2=10000"

"P=100"

Therefore, the market clearing Prices is 100

"Q^s= 30\times 100=3000"

"=3000"

Therefore market clearing Quantity is 3000


Sketch